OpenStudy (anonymous):

Which equation is related to this equation? k ÷ 4 = 13 A. k = 13 – 4 B. k = 13 ÷ 4 C. k = 13 ∙ 4 D. k = 4 ÷ 13

OpenStudy (anonymous):

help

OpenStudy (leonardo0430):

gtg hope i helped, read mail

OpenStudy (anonymous):

who evr helps do you mind helping me with a few other

OpenStudy (leonardo0430):

i gtg :/ cant stay

OpenStudy (anonymous):

oh can you refer someone here

OpenStudy (leonardo0430):

@PRAETORIAN.10 help her i gtg and no i dont care if your a doctor you still had math to do in highschool right?

OpenStudy (leonardo0430):

@jane11509 help her :l i gtg

OpenStudy (leonardo0430):

rest of my friend are offline :(

OpenStudy (anonymous):

oh well thankz for your help

OpenStudy (leonardo0430):

no problem if you got any questions mail me back :P

OpenStudy (anonymous):

k will do

OpenStudy (anonymous):

Its C

OpenStudy (leonardo0430):

YES PEOPLE HELP HER I GTG :L

OpenStudy (leonardo0430):

bye ^_^

OpenStudy (anonymous):

bye

OpenStudy (leonardo0430):

tell the your other question.........

OpenStudy (anonymous):

Pretty Sure it is C Koolklue22 did that help?

OpenStudy (leonardo0430):

yes its C

OpenStudy (anonymous):

@fatima can you help me with a couple more

OpenStudy (anonymous):

Sure-

OpenStudy (anonymous):

Which equation is related to the given equation? Column A Column B 1. 61 + 18 = z 2. z = 61 – 18 3. –18 = 61 – z 4. 18 = 61 – z A. z + 18 = 61 B. z – 18 = 61

OpenStudy (anonymous):

Sorry- that is really confusing- can you paste a screen shot of whatever q ur doing? The table isnt coming through well

OpenStudy (anonymous):

oh i copy pasted hold on

OpenStudy (anonymous):

Use a related equation that has the variable on one side. Then simplify the other side. = 16 A. m = 64 B. m = 4 C. m = 12 D. m = 20

OpenStudy (anonymous):

its 

OpenStudy (anonymous):

m^2= 16, so I am guessing that m=4. The __=16 is throwing me off

OpenStudy (anonymous):

thats the missing part

OpenStudy (anonymous):

yeah- its definitely m=4

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

any more?

OpenStudy (anonymous):

. Use a related equation that has the variable on one side. Then simplify the other side. h – 8 = 32 A. B. C. D.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Can you screen shot again? the ◘s are not very...good

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

h – 8 = 32 A. B. C. D.

OpenStudy (anonymous):

srry hold on

OpenStudy (anonymous):